# The PAU Survey: galaxy stellar population properties estimate with narrow-band data

### v1

@@
11institutetext: Institute for Astro- and Particle Physics, Universität Innsbruck, Technikerstraße 25/8, 6020 Innsbruck, Austria 22institutetext: University Observatory, Faculty of Physics, Ludwig-Maximilians-Universität, Scheinerstr. 1, 81677 Munich, Germany 33institutetext: Excellence Cluster ORIGINS, Boltzmannstr. 2, 85748 Garching, Germany 44institutetext: Institute of Space Sciences (ICE, CSIC), Campus UAB, Carrerde Can Magrans, 08193 Bellaterra (Barcelona), Spain 55institutetext: Instituto Astrofísica de Canarias, Av. Via Láctea s/n, E38205 La Laguna, Spain 66institutetext: Department of Astronomy, Tsinghua University, Beijing 100084, China 77institutetext: Centro de Investigaciones Energéticas, Medioambientales y Tecnológicas (CIEMAT), Avenida Complutense 40, E-28040 Madrid (Madrid), Spain 88institutetext: Port d’Informació Científica (PIC), Campus UAB, C. Albareda s/n, 08193 Bellaterra (Cerdanyola del Vallès), Spain 99institutetext: Institució Catalana de Recerca i Estudis Avançats (ICREA), Pg. de Lluís Companys 23, 08010 Barcelona, Spain 1010institutetext: Institut de Física d’Altes Energies (IFAE), The Barcelona Institute of Science and Technology, Campus UAB, 08193 Bellaterra (Barcelona) Spain $1$institutetext: Instituto de Fisica Teorica (UAM/CSIC), Universidad Autonoma de Madrid, E-28049 Madrid, Spain 1212institutetext: Institute of Cosmology and Gravitation (University of Portsmouth), Portsmouth, UK 1313institutetext: Institut d’Estudis Espacials de Catalunya (IEEC), Gran Capità 2-4, Ed. Nexus 201, 08034 Barcelona, Spain 1414institutetext: Satlantis, University Science Park, Sede Bld 48940, Leioa-Bilbao, Spain 1515institutetext: Jockey Club Institute for Advanced Study, The Hong Kong University of Science and Technology, Hong Kong S.A.R., China 1616institutetext: Institute for Computational Cosmology, Department of Physics, Durham University, South Road, Durham DH1 3LE, UK
The PAU Survey: galaxy stellar population properties estimate with narrow-band data
@@
Narrow-band galaxy surveys have recently gained interest as a promising method to achieve the necessary accuracy on the photometric redshift estimate of individual galaxies for next-generation stage-IV cosmological surveys. One key advantage is the ability to provide higher spectral resolution information about galaxies that should allow a more accurate and precise estimation of galaxy stellar population properties. However, the impact of adding narrow-band photometry on the stellar population properties estimate is largely unexplored. The scope of this work is two-fold: on one side, leveraging the predictive power of broad-band and narrow-band data to infer galaxy physical properties such as stellar masses, ages, star formation rates and metallicities. On the other hand, evaluating the improvement of performance in estimating galaxy properties when we use narrow-band data instead of broad-band. In this work we measure the stellar population properties of a sample of galaxies in the COSMOS field for which both narrow-band and broad-band data are available. In particular, we employ narrow-band data from the Physics of the Accelerating Universe Survey (PAUS) and broad-band data from the Canada France Hawaii Telescope legacy survey (CFHTLS). We use two different spectral energy distribution fitting codes to measure galaxy properties, namely cigale and Prospector. We find that the increased spectral resolution of narrow-band photometry does not yield a substantial improvement on constraining galaxy properties using spectral energy distribution fitting. Still we find that we obtain a more diverse distribution of metallicities and dust optical depths with cigale when employing the narrow-band data. The effect is not as prominent as expected, which we relate this to the low narrow-band signal-to-noise ratio of a majority of the galaxies, the respective drawbacks of both codes as well as the coverage only in the optical regime. The measured properties are afterwards compared to the ones in the COSMOS2020 catalogue, showing good agreement. We release the catalogue of measured properties alongside this work.
@@ 1 Introduction
The stellar content of galaxy populations offers important insights into galaxy evolution and formation, as the physical properties can provide a direct indicator of evolutionary paths to trace star formation histories and metal enrichment. At the same time, the physical properties of galaxies are predictive of effects relevant for cosmological studies that utilize galaxies, particularly their clustering bias and the degree of the intrinsic alignment of a galaxy’s shape relative to the local tidal field (Benson et al., 2000; Li et al., 2006; Joachimi et al., 2013; Samuroff et al., 2021). It is therefore vital to determine the distribution and the evolution of stellar populations to constrain models of galaxy evolution and answer current cosmological questions.
@@
The most used technique to obtain galaxy stellar population properties from photometric data is spectral energy distribution (SED) fitting (for a review see e.g. Conroy, 2013). This process requires mainly three ingredients, namely the observational data itself, a physical model of the galaxy population, and a method of statistical inference to connect them. The fluxes or apparent magnitudes of observed or simulated galaxies are related not only to the underlying stellar population, but also depend on redshift, the emission from gas and dust in the interstellar medium (ISM), the history of star formation and the dust attenuation of light along the line of sight. A template-based fitting process hence has to account for these contributions, which has the potential to introduce a large number of degrees of freedom into the routine, increasing consequently the complexity of the estimation.
@@
Optimization of the SED fitting parameter space and a statistically powerful and efficient approach is vital for the characterization of galaxy populations in large-scale surveys. Therefore a collection of panchromatic SED codes have been developed that use stellar population synthesis (SPS) to forward-model the galaxy population and fit a spectrum motivated by the physical processes that account for the observed spectral distributions (Walcher et al., 2011). These codes employ a variety of different frameworks or templates as well as various Bayesian methods, such as Markov-Chain Monte Carlo sampling or $\chi^{2}$ minimization on a model grid (Conroy, 2013; Thorne et al., 2021; Pacifici et al., 2023).
@@
SED fitting codes work by fitting real galaxy fluxes with ones generated from template SEDs. The higher the resolution of the photometric data, the more reliable will be the galaxy properties estimated via SED fitting. Additionally, the characteristics of the models generated by the fitting code influence the quality of the estimation, depending on their ability to describe the plethora of observed galaxies accurately. Different initial mass functions (IMFs) or dust attenuation laws for instance, as well as assumptions about the stellar metallicity, may introduce a difference in stellar mass estimates at a $\sim 0.2$ dex level (see e.g. Speagle et al., 2014; Bernardi et al., 2018) or bimodality in the stellar mass distribution (Mitchell et al., 2013). Previous tests on mock catalogs by Mejía-Narváez et al. (2017) determined an improved accuracy in physical parameter estimation from the Javalambre-PAU Astrophysical Survey (J-PAS) narrow-band data in comparison to broad-band fluxes for passive galaxies only. Additionally, studies with PAUS mocks showed high capability to recover galaxy color-redshift relations (Manzoni et al., 2024).
@@
Narrow-band surveys can be a powerful niche between spectroscopy and photometry for this task. While wavelength-resolved spectroscopic data only comes at the detriment of target selection biases, flux-limitation and long integration times, broad-band photometric surveys lack spectral resolution. An increased number of filter bands with narrow bandpasses however can deliver a quasi-spectroscopic energy resolution with high accuracy photometric redshifts, while the imaging approach ensures time efficiency and thus a large sample size, which is of paramount importance for statistical cosmological research and the creation of a global model of the galaxy population. One such narrow-band survey is the Physics of the Accelerating Universe Survey (hereafter, PAUS) conducted at the Obervatorio del Roque de los Muchachos on La Palma, Spain (Benítez et al., 2009; Martí et al., 2014; Serrano et al., 2023). The survey uses 40 uniformly spaced narrow-band filters with a full width at half maximum (FWHM) of 12.5 nm in the optical and achieves photo-$z$ precisions up to $\sigma(z)/(1+z)\sim 0.0057$ for $i_{\text{AB}}<22.5$ over a volume of 0.3 (GPc/h)3 (Eriksen et al., 2019; Alarcon et al., 2021; Soo et al., 2021; Cabayol et al., 2023; Navarro-Gironés et al., 2023).
@@
In this work we employ narrow-band data from PAUS and broad-band data from the Canada-France-Hawaii Telescope Legacy Survey (CFHTLS, Cuillandre et al., 2012; Gwyn, 2012) to determine galaxy stellar population properties in the PAUS-COSMOS field with two SED fitting codes, namely cigale (Burgarella et al., 2005; Noll et al., 2009; Boquien et al., 2019) and Prospector (Leja et al., 2017; Johnson et al., 2021). We apply the SED fitting routine to a sample of PAUS-observed galaxies, fitting both narrow-bands and broad-bands combined and broad-bands data alone, to determine the distribution of properties, quantify the benefit of the PAUS narrow-band filter set and identify drawbacks and advantages of the two codes.
The two SED fitting codes have a range of hyper-parameters that can be tuned by the user. Therefore, in order to identify the best set of hyper-parameters that can be later applied to data, we generate a set of synthetic PAUS-like galaxies simulated with Prospector and the underlying FSPS (Flexible Stellar Population Synthesis) framework (Conroy et al., 2009; Conroy & Gunn, 2010; Johnson et al., 2022). The knowledge of the true values of simulated galaxy properties allows us to adjust the hyper-parameter space and benchmark the two codes against one another. The tuned parameter space is then used to measure stellar masses, star formation rates (SFRs) and mass-weighted ages for observed galaxies in the COSMOS field. The results are then compared to the COSMOS2020 catalog (Weaver et al., 2022) to which our sample is cross-matched by RA and DEC.
Section 2 describes the PAUS and CFHTLS observational data used in our work. Section 3 presents in detail the two SED fitting codes cigale and Prospector. Furthermore, we also provide a description of how we generate the library of mock galaxies and the priors. In Sect. 4 we tune the hyper-parameter of cigale and Prospector by fitting mock galaxy photometry and compare the results between the two codes. Section 5 uses the tuned hyper-parameter space to measure the stellar population properties of observed galaxies in the COSMOS field with narrow-band+broad-bands, as well as broad-bands only, and compares them against COSMOS2020 catalogue properties.
Throughout this work, we assume a flat $\Lambda$CDM cosmology with $\Omega_{m}=0.32$, $\Omega_{\Lambda}=0.68$, $H_{0}=67.4\,\mathrm{km\,s}^{-1}\,\mathrm{Mpc}^{-1}$ (Planck Collaboration et al., 2020).
@@ 2 Observational Data
This section describes the observational data analyzed in this work. The PAUS filter set contains 40 equally spaced slightly overlapping narrow-band filters with a FWHM of 12.5 Å in the spectral range from 4550 Å to 8450 Å. The total throughputs of the filters, including atmospheric effects as well as telescope, filter and detector influences, like the CCD quantum efficiency, are shown in Fig. 1. PAUCam has 18 red-sensitive full depletion CCDs, of which only the 8 central ones are used for narrow-band imaging due to vignetting and image curvature. Its filters were commissioned to the prime focus of the William Herschel Telescope and they cover a total FOV of $\sim$1 deg2 at an average pixel scale of 0.265” (Benítez et al., 2009; Castander et al., 2012). We use fluxes, flux errors and and photometric redshifts from the PAUS+COSMOS v0.4 flux/photo-$z$ catalogs hosted on CosmoHub (Alarcon et al., 2021).
The additional broad-band photometry is taken from the final data release of CFHTLS (Cuillandre et al., 2012). This survey has been conducted at Mauna Kea, Hawaii, using MegaCam (Boulade et al., 2003) and covers 175 deg2 in $ugriz$ bands with sub-arcsecond seeing at a pixel scale of 0.187”. Figure 1 shows a plot of the total throughput of PAUS narrow-bands and the complementary CFHTLS broad-band filters.
Both these surveys have common coverage in the COSMOS field ($\hat{=}$ CFHTLS D2 field), which is one of the most widely observed fields in astrophysics. Originally investigated by the Cosmic Evolution Survey (COSMOS, Scoville et al., 2007) there exists today multi-wavelength data over a $\sim$2 deg2 surface centred at RA = 150.12 deg (10:00:28.6) DEC = +2.21 deg (+02:12:21.0) (J2000) near the celestial equator. More recent COSMOS catalogues of improved photometric redshifts and physical property estimates using a broader wavelength baseline including narrow- and medium-band filters have since then been released (Laigle et al., 2016; Weaver et al., 2022), the latter of which is used in this work for comparison to our results. We perform the analysis on the common sample of both surveys, consisting of 25,491 galaxies up to redshift $z\sim 2$ and $i_{\mathrm{AB}}\sim 23.5$, all of which have observations in all 40 PAUS bands and the five CFHTLS $ugriz$ bands.
Figure 1: Total throughput (including the effects of atmosphere, telescope optics, filter design and detector assembly) of the 40 PAUS narrow-band filters (colored) spanning the wavelength range from 4550 Å to 8450 Å and the CFHTLS $ugriz$ broad-band filters (gray).

### v2

# The PAU Survey: Galaxy stellar population properties estimates with narrowband data

@@
11institutetext: Universität Innsbruck, Institut für Astro- und Teilchenphysik, Technikerstr. 25/8, 6020 Innsbruck, Austria 22institutetext: University Observatory, Faculty of Physics, Ludwig-Maximilians-Universität, Scheinerstr. 1, 81677 Munich, Germany 33institutetext: Excellence Cluster ORIGINS, Boltzmannstr. 2, 85748 Garching, Germany 44institutetext: Institute of Space Sciences (ICE, CSIC), Campus UAB, Carrerde Can Magrans, 08193 Bellaterra (Barcelona), Spain 55institutetext: Instituto Astrofísica de Canarias, Av. Via Láctea s/n, E38205 La Laguna, Spain 66institutetext: Department of Astronomy, Tsinghua University, Beijing 100084, China 77institutetext: Centro de Investigaciones Energéticas, Medioambientales y Tecnológicas (CIEMAT), Avenida Complutense 40, E-28040 Madrid (Madrid), Spain 88institutetext: Port d’Informació Científica (PIC), Campus UAB, C. Albareda s/n, 08193 Bellaterra (Cerdanyola del Vallès), Spain 99institutetext: Institució Catalana de Recerca i Estudis Avançats (ICREA), Pg. de Lluís Companys 23, 08010 Barcelona, Spain 1010institutetext: Institut de Física d’Altes Energies (IFAE), The Barcelona Institute of Science and Technology, Campus UAB, 08193 Bellaterra (Barcelona) Spain $1$institutetext: Instituto de Fisica Teorica (UAM/CSIC), Universidad Autonoma de Madrid, E-28049 Madrid, Spain 1212institutetext: Institute of Cosmology and Gravitation (University of Portsmouth), Portsmouth, UK 1313institutetext: Institut d’Estudis Espacials de Catalunya (IEEC), Gran Capità 2-4, Ed. Nexus 201, 08034 Barcelona, Spain 1414institutetext: Satlantis, University Science Park, Sede Bld 48940, Leioa-Bilbao, Spain 1515institutetext: Jockey Club Institute for Advanced Study, The Hong Kong University of Science and Technology, Hong Kong S.A.R., China 1616institutetext: Institute for Computational Cosmology, Department of Physics, Durham University, South Road, Durham DH1 3LE, UK
The PAU Survey: Galaxy stellar population properties estimates with narrowband data
@@
A newfound interest has been seen in narrowband galaxy surveys as a promising method for achieving the necessary accuracy on the photometric redshift estimate of individual galaxies for next-generation stage IV cosmological surveys. One key advantage is the ability to provide higher spectral resolution information on galaxies, which ought to allow for a more accurate and precise estimation of the stellar population properties for galaxies. However, the impact of adding narrowband photometry on the stellar population properties estimate is largely unexplored. The scope of this work is two-fold: 1) we leverage the predictive power of broadband and narrowband data to infer galaxy physical properties, such as stellar masses, ages, star formation rates, and metallicities; and 2) we evaluate the improvement of performance in estimating galaxy properties when we use narrowband instead of broadband data. In this work, we measured the stellar population properties of a sample of galaxies in the COSMOS field for which both narrowband and broadband data are available. In particular, we employed narrowband data from the Physics of the Accelerating Universe Survey (PAUS) and broadband data from the Canada France Hawaii Telescope legacy survey (CFHTLS). We used two different spectral energy distribution (SED) fitting codes to measure galaxy properties, namely, cigale and Prospector. We find that the increased spectral resolution of narrowband photometry does not yield a substantial improvement in terms of constraining the galaxy properties using the SED fitting. Nonetheless, we find that we are able to obtain a more diverse distribution of metallicities and dust optical depths with cigale when employing the narrowband data. The effect is not as prominent as expected, which we relate to the low narrowband signal-to-noise ratio (S/N) of a majority of the sampled galaxies, the respective drawbacks of both codes, and the restriction of coverage to the optical regime. The measured properties are compared to those reported in the COSMOS2020 catalogue, showing a good agreement. We have released the catalogue of measured properties in tandem with this work.
@@ 1 Introduction
The stellar content of galaxy populations offers important insights into galaxy evolution and formation, as the physical properties can provide a direct indicator of evolutionary paths for tracing star formation histories and metal enrichment. At the same time, the physical properties of galaxies are predictive with respect to effects relevant to cosmological studies of galaxies, particularly for their clustering bias and the degree of the intrinsic alignment of a galaxy’s shape relative to the local tidal field (Benson et al., 2000; Li et al., 2006; Joachimi et al., 2013; Samuroff et al., 2021). It is therefore vital to determine the distribution and the evolution of stellar populations to constrain models of galaxy evolution and answer current cosmological questions.
@@
The common technique used to obtain galaxy stellar population properties from photometric data is spectral energy distribution (SED) fitting (for a review, see e.g., Conroy, 2013). This process requires three main ingredients, namely: the observational data, a physical model of the galaxy population, and a method of statistical inference to connect them. The fluxes or apparent magnitudes of observed or simulated galaxies are related not only to the underlying stellar population, but also depend on redshift, the emission from gas and dust in the interstellar medium (ISM), the history of star formation, and the dust attenuation of light along the line of sight. A template-based fitting process hence has to account for these contributions, which has the potential to introduce a large number of degrees of freedom into the routine, increasing consequently the complexity of the estimation.
@@
The optimization of the SED fitting parameter space and a statistically powerful and efficient approach is vital for the characterization of galaxy populations in large-scale surveys. Therefore, a collection of panchromatic SED codes have been developed based on stellar population synthesis (SPS) to forward-model the galaxy population and fit a spectrum motivated by the physical processes that account for the observed spectral distributions (Walcher et al., 2011). These codes employ a variety of different frameworks or templates, as well as various Bayesian methods, such as Markov chain Monte Carlo sampling or $\chi^{2}$ minimization on a model grid (Conroy, 2013; Thorne et al., 2021; Pacifici et al., 2023).
@@
These SED fitting codes work by fitting real galaxy fluxes with ones generated from template SEDs. The higher the resolution of the photometric data, the more reliable will be the galaxy properties estimated via SED fitting. Additionally, the characteristics of the models generated by the fitting code influence the quality of the estimation, depending on their ability to describe the plethora of observed galaxies accurately. Different initial mass functions (IMFs) or dust attenuation laws for instance, as well as assumptions about the stellar metallicity, may introduce a difference in stellar mass estimates at a $\sim 0.2$ dex level (see e.g., Speagle et al., 2014; Bernardi et al., 2018) or bimodality in the stellar mass distribution (Mitchell et al., 2013). Previous tests on mock catalogs by Mejía-Narváez et al. (2017) determined an improved accuracy in physical parameter estimation from the Javalambre-PAU Astrophysical Survey (J-PAS) narrowband data in comparison to broadband fluxes for passive galaxies only. Additionally, studies with PAUS mocks showed high capability to recover galaxy color-redshift relations (Manzoni et al., 2024).
@@
Narrowband surveys can be a powerful niche between spectroscopy and photometry for this task. While wavelength-resolved spectroscopic data only comes at a detriment to target selection biases, flux-limitation, and long integration times, broadband photometric surveys lack spectral resolution. However, an increased number of filter bands with narrow bandpasses can deliver a quasi-spectroscopic energy resolution with high accuracy photometric redshifts, while the imaging approach ensures time efficiency and thus a large sample size, which are of paramount importance for statistical cosmological research and the creation of a global model of the galaxy population. One such narrowband survey is the Physics of the Accelerating Universe Survey (PAUS) conducted at the Obervatorio del Roque de los Muchachos on La Palma, Spain (Benítez et al., 2009; Martí et al., 2014; Serrano et al., 2023). The survey uses 40 uniformly spaced narrowband filters with a full width at half maximum (FWHM) of 12.5 nm in the optical and achieves photo-$z$ precisions up to $\sigma(z)/(1+z)\sim 0.0057$ for $i_{\text{AB}}<22.5$ over a volume of 0.3 (GPc/h)3 (Eriksen et al., 2019; Alarcon et al., 2021; Soo et al., 2021; Cabayol et al., 2023; Navarro-Gironés et al., 2023).
@@
In this work, we employ narrowband data from PAUS and broadband data from the Canada-France-Hawaii Telescope Legacy Survey (CFHTLS, Cuillandre et al., 2012; Gwyn, 2012) to determine galaxy stellar population properties in the PAUS-COSMOS field with two SED fitting codes, namely cigale (Burgarella et al., 2005; Noll et al., 2009; Boquien et al., 2019) and Prospector (Leja et al., 2017; Johnson et al., 2021). We apply the SED fitting routine to a sample of PAUS-observed galaxies, fitting both the narrowband and broadband combined as well as the broadband data alone, to determine the distribution of properties, quantify the benefit of the PAUS narrowband filter set, and identify drawbacks and advantages of the two codes.
The two SED fitting codes have a range of hyper-parameters that can be tuned by the user. Therefore, in order to identify the best set of hyper-parameters that can be later applied to data, we generated a set of synthetic PAUS-like galaxies simulated with Prospector and the underlying Flexible Stellar Population Synthesis (FSPS) framework (Conroy et al., 2009; Conroy & Gunn, 2010; Johnson et al., 2022). The knowledge of the true values of simulated galaxy properties allows us to adjust the hyper-parameter space and benchmark the two codes against one another. The tuned parameter space is then used to measure stellar masses, star formation rates (SFRs) and mass-weighted ages for observed galaxies in the COSMOS field. The results are then compared to the COSMOS2020 catalog (Weaver et al., 2022), to which our sample is cross-matched by RA and DEC.
Section 2 describes the PAUS and CFHTLS observational data used in our work. Section 3 presents in detail the two SED fitting codes cigale and Prospector. Furthermore, we also provide a description of how we generate the library of mock galaxies and the priors. In Sect. 4, we tune the hyper-parameter of cigale and Prospector by fitting mock galaxy photometry and compare the results between the two codes. Section 5 uses the tuned hyper-parameter space to measure the stellar population properties of observed galaxies in the COSMOS field with narrowband+broadband, as well as broadband only, and compares them against COSMOS2020 catalogue properties. Throughout this work, we assume a flat $\Lambda$CDM cosmology with $\Omega_{m}=0.32$, $\Omega_{\Lambda}=0.68$, and $H_{0}=67.4\,\mathrm{km\,s}^{-1}\,\mathrm{Mpc}^{-1}$ (Planck Collaboration et al., 2020).
@@ 2 Observational data
This section describes the observational data analyzed in this work. The PAUS filter set contains 40 equally spaced slightly overlapping narrowband filters with a FWHM of 12.5 Å in the spectral range from 4550 Å to 8450 Å. The total throughputs of the filters, including atmospheric effects, as well as telescope, filter, and detector effects, such as the CCD quantum efficiency, are shown in Fig. 1. PAUCam has 18 red-sensitive full depletion CCDs, of which only the 8 central ones are used for narrowband imaging due to vignetting and image curvature. Its filters were commissioned to the prime focus of the William Herschel Telescope and they cover a total FOV of $\sim$1 deg2 at an average pixel scale of 0.265” (Benítez et al., 2009; Castander et al., 2012). We use fluxes, flux errors and and photometric redshifts from the PAUS+COSMOS v0.4 flux/photo-$z$ catalogs hosted on CosmoHub (Alarcon et al., 2021).
The additional broadband photometry was taken from the final data release of CFHTLS (Cuillandre et al., 2012). This survey has been conducted at Mauna Kea, Hawaii, using MegaCam (Boulade et al., 2003) and covers 175 deg2 in $ugriz$ bands with sub-arcsecond seeing at a pixel scale of 0.187”. Figure 1 shows a plot of the total throughput of PAUS narrowbands and the complementary CFHTLS broadband filters.
Both these surveys have common coverage in the COSMOS field ($\hat{=}$ CFHTLS D2 field), which is one of the most widely observed fields in astrophysics. Originally investigated by the Cosmic Evolution Survey (COSMOS, Scoville et al., 2007) today there are multi-wavelength data available over a $\sim$2 deg2 surface centred at RA = 150.12 deg (10:00:28.6) DEC = +2.21 deg (+02:12:21.0) (J2000) near the celestial equator. More recent COSMOS catalogues of improved photometric redshifts and physical property estimates using a broader wavelength baseline including narrow- and medium band filters have since been released (Laigle et al., 2016; Weaver et al., 2022), the latter of which is used in this work for comparison to our results. We performed the analysis on the common sample of both surveys, consisting of 25,491 galaxies up to redshift $z\sim 2$ and $i_{\mathrm{AB}}\sim 23.5$, all of which have observations in all 40 PAUS bands and the five CFHTLS $ugriz$ bands.
Figure 1: Total throughput (including the effects of atmosphere, telescope optics, filter design and detector assembly) of the 40 PAUS narrowband filters (colored) spanning the wavelength range from 4550 Å to 8450 Å and the CFHTLS $ugriz$ broadband filters (gray).